LIF AND MICROWAVE SPECTROSCOPY OF CH_{2}CHS

\underline{MASAKAZU NAKAJIMA}, AKIRA MIYOSHI, Department of Chemical System Engineering, The University of Tokyo, Tokyo, 113-8656, Japan; YOSHIHIRO SUMIYOSHI, and YASUKI ENDO, Department of Basic Science, The University of Tokyo, Tokyo, 153-8902, Japan.

Precise molecular constants of the CH_{2}CHS (vinylthio) radical in the ground vibronic level were determined by Fouriertransform microwave (FTMW) and FTMW-millimeter wave double-resonance spectroscopy^{a}. The B̃ – X̃ electronic transition of CH_{2}CHS was observed by LIF spectroscopy. Rotational constants in the upper electronic state were determined from a rotationally resolved LIF excitation spectrum of jet-cooled CH_{2}CHS. A dispersed fluorescence spectrum from the zero-vibtational level of the B̃ state was also measured to determine vibrational frequencies in the ground electronic state. The experimental molecular constants were compared with results of ab initio calculations.

aY. Sumiyoshi et al., J. Chem. Phys. \underline{123}, 054324 (2005).